Solve the differential equation

xcos(y/x) (ydx+xdy) = ysin(y/x) (xdy-ydx)


1
Expert's answer
2021-06-06T16:37:38-0400
"x\cos(\dfrac{y}{x})(ydx+xdy)=y\sin(\dfrac{y}{x})(xdy-ydx)"

"\dfrac{ydx+xdy}{xy}=\tan(\dfrac{y}{x})\dfrac{xdy-ydx}{x^2}"


"\dfrac{d(xy)}{xy}=\tan(\dfrac{y}{x})d(\dfrac{y}{x})"

"\int\dfrac{d(xy)}{xy}=\int\tan(\dfrac{y}{x})d(\dfrac{y}{x})"

"\int \tan u du=\int\dfrac{\sin u}{\cos u}du"

"=-\ln|\cos u|+C_1=\ln|\sec u|+c_1"

"\ln|xy|=\ln|\sec (\dfrac{y}{x})|-\ln c"


"\sec (\dfrac{y}{x})=cxy"